# linecolor=white, backgroundcolor=gray!20

### v1

@@ 4.2 Parameterization with Adapter-guided Protein Language Models
To efficiently learn $(a^{i},\boldsymbol{x}^{i},O^{i},\boldsymbol{\chi}^{i})$ for every residue, we adopt and improve the FramePred architecture introduced in FrameDiff [80] and FrameFlow [79]. It effectively incorporates the Invariant Point Attention (IPA) [32] to encode spatial features and ensure the equivariance. IPA consists of multiple Transformer layers [68] to capture sequence-level attributes. In addition, considering the periodicity, torsion angles $\boldsymbol{\chi}\in[0,2\pi)^{7}$ are flexibly encoded by applying multi-frequency sine and cosine transformations [41], which are also fused with the timestep embedding and residue sequence embedding into IPAs.
PLMs capture the evolutionary patterns from large-scale sequence data, and this knowledge is proven to be supportive of protein folding [43] and inverse design [84], which is evidently beneficial for our co-design task as well. Some prior studies [77] integrate this knowledge by immediately appending the geometric networks to PLMs. Drawing inspiration from LM-Design [84], we employ a lightweight structural adapter to endow PLMs with structural awareness. In our approach, if the output of IPA is denoted as $\boldsymbol{h}_{\text{IPA}}\in\mathbb{R}^{\psi_{\text{IPA}}}$, the attention at the $l$-th layer is computed as: $\boldsymbol{o}=\mathrm{softmax}\left(\frac{\boldsymbol{h}_{\text{seq}}\boldsymbol{W}_{Q}\cdot\boldsymbol{h}_{\text{IPA}}^{\top}\boldsymbol{W}_{V}^{\top}}{\sqrt{\psi_{\text{seq}}}}\right)\boldsymbol{h}_{\text{IPA}}\boldsymbol{W}_{V}$, where $\boldsymbol{h}_{\text{seq}}\in\mathbb{R}^{\psi_{\text{seq}}}$ is the sequential embedding of the last $(l-1)$-th layer. $\boldsymbol{W}_{K}$, $\boldsymbol{W}_{Q}$, and $\boldsymbol{W}_{V}$ are trainable weights for key, query, and values, separately. This module allows the usage of both sequence-based and structure-based information for protein understanding.
@@ 4.3 Controlling Flow Matching with Target Proteins
Existing de novo design algorithms [79, 80] are mainly targeting monomers or biomolecules without receptors. Some recent studies [46, 41, 42] have considered binder design but their performance is strictly restricted by the number of available complex structures (e.g., SabDAB [20] and PepMerge [41]). To bridge the gap and exploit all crystal structures, we propose to pretrain our flow model on the vast amount of general proteins (i.e., Protein Data Bank) and then fine-tune it on the target-specific design challenges.
To begin with, $\mathcal{F}_{\Theta}:\mathcal{P}_{0}\rightarrow\mathcal{C}_{\text{pro}}$ is a co-design flow model without target awareness, which can de novo design a protein ${C}_{\text{pro}}$ from any initial protein ${C}_{0}$ that is drawn from a prior distribution $p_{0}$. It is trained on general proteins to approximate $p(a,\boldsymbol{x},O,\boldsymbol{\chi})$. To enable the awareness of $\mathcal{F}_{\Theta}$ to the receptor $C_{\text{rec}}$, we borrow ideology from the popular ControlNet architecture [82], locking (freeze) the parameters $\Theta$ of the original block and simultaneously cloning the block to a trainable copy with parameters $\Theta_{c}$. The trainable copy takes a receptor protein $C_{\text{rec}}$ as input and is connected to the locked model with zero IPAs denoted as $\mathcal{Z}(\cdot)$. Notably, both weight and bias in $\mathcal{Z}(\cdot)$ are initialized to zeros. In practice, two instances of zero IPAs are used with parameters $\Theta_{\mathrm{z}1}$ and $\Theta_{\mathrm{z}2}$, respectively:
@@
Table 1: Evaluation of methods in the traditional L-peptide sequence-structure co-design task. The best and suboptimal results are labeled boldly and underlined, respectively. Baselines with $\dagger$ are reproduced.
@@
[…]
@@ 4.4 D-peptide Design
In this work, we generate mini D-protein binders, specifically D-peptides, using the mirror-image algorithm [63, 54]. The mirror-image operation, denoted as $\Psi(\cdot)$, transforms any given protein structure into its corresponding mirror image. $\Psi(\cdot)$ inverts the spatial configuration of the protein, effectively reflecting it across an imaginary plane while preserving the relative distances and angles between atoms. In our approach, we first apply $\Psi(\cdot)$ to the receptor protein $C_{\text{rec}}$, converting it into its mirror-image counterpart. Then, we pass this mirrored receptor through a flow model $f_{\theta}(\cdot):\mathcal{C}_{\text{rec}}\rightarrow\mathcal{C}_{\text{pep}}$, which is trained on natural-chirality (L-) proteins, to yield the corresponding peptide. Finally, we apply the inverse operation $\Psi^{-1}(\cdot)$ to convert the resulting peptide back to the D-configuration. Formally, the generation process can be expressed as: {mdframed}
$C^{\text{D-pep}}=\Psi^{-1}(f_{\theta}(\Psi(C_{\text{rec}}))).$ (20)
5 Experiments
We execute two types of experiments to validate the effectiveness of our D-Flow approach. The first, detailed in Sec. 5.1, involves the conventional co-design challenge for L-proteins, where models generate peptides’ both sequences and structures based on a given receptor binding site. The second experiment, described in Sec. 5.2, focuses on generating D-peptides. For benchmarking, we use the PepMerge dataset [41], derived from PepBDB [73] and Q-BioLip [72]. To ensure a fair comparison with prior work [41], we cluster peptide-protein complexes based on 40% sequence identity using MMseqs2 [61], after removing duplicates and applying empirical filters (e.g., resolution $<$ 4Å, peptide length between 3 and 25). This process results in 8,365 non-redundant complexes across 292 clusters. For consistency in comparison, we use the same test set as Li et al. [41], consisting of 10 clusters and 158 complexes. Pretraining data contain monomers between lengths 60 and 384 with resolution $<5\textup{\AA}$ downloaded from PDB [8] on August 8, 2021, ensuring no data leakage for peptides. The data is then filtered by including proteins with high secondary structure compositions only. Monomers with more than 50% loops are also removed using DSSP [33], resulting in 18,684 proteins.
5.1 Unconditioned Sequence-structure Co-design
Baselines.
We select two categories of state-of-the-art protein design methods as baselines. The first category disregards side-chain conformations and includes approaches such as RFDiffusion [71] and ProteinGen [45]. RFDiffusion generates protein backbones, with sequences predicted afterward using ProteinMPNN [15], while ProteinGen improves on RFDiffusion by jointly sampling both backbones and sequences. The second category considers full-atom protein generation, containing Diffusion [46], PPIFlow [42], and PepFlow [41]. PepFlow has three variants based on whether backbones, sequences, and side-chain angles are sampled.
Metrics.
Generated peptides are evaluated across three key aspects. (1) Geometry: Designed peptides should closely mirror native sequences and structures. We quantify sequence similarity through the amino acid recovery rate (AAR), which measures sequence identity between generated and ground truth peptides. Structural similarity is evaluated using the root-mean-square deviation (RMSD) of $C_{\alpha}$ atoms after complex alignment. The secondary-structure similarity ratio (SSR) captures the proportion of matching secondary structures, while the binding site ratio (BSR) evaluates the overlap between generated and native peptide binding sites on the target protein. (2) Energy: We aim to design high-affinity peptide binders that enhance protein-peptide complex stability. Affinity represents the percentage of generated peptides achieving higher binding affinities (lower binding energies) than the native peptide. Stability indicates the fraction of complexes exhibiting lower total energy than the native state. All energy calculations are performed using Rosetta [3]. (3) Design: Designability measures the consistency between designed sequences and structures, calculated as the proportion of sequences that fold into structures similar to their generated conformations ($C_{\alpha}$ RMSD $<$ 2 Å). Sequence refolding is performed using ESMFold [43]. Diversity, computed as the average of one minus the pairwise TM-Score [83], quantifies structural variation among the designed peptides.
Main Results.
As documented in Tab. 1, D-Flow generates peptide sequences with the closest resemblance to native ones, achieving a high amino acid identity of 58.69%, a 14.51% improvement over PepFlow. It excels in all geometry-related metrics, highlighting its close alignment with the binding sites of native peptides. Besides, D-Flow scores 75.14% in designability and 0.60 in diversity, maintaining a good balance between structural fidelity and variety. Moreover, D-Flow also demonstrates superior energy-based properties, achieving the best stability score at 26.85% and the best affinity score at 24.31%, which are critical for the formation of strong and stable complexes. In a nutshell, D-Flow outpasses all baselines across most key metrics, indicating its strength in consistently producing the most accurate peptide sequences and structures along with the optimal stability and affinity for their targets.
We also navigate the contribution of different components of D-Flow, containing the discrete flow matching (DFM), structural-adapted PLM, the pretraining on PDB, and the ControlNet-style transfer learning technique. Tab. 2 reports the ablation study results. It can be observed that the integration of additional unlabeled structural data yielded the most substantial improvement, increasing AAR by 5.93%. This supports our hypothesis that limited binding complex data significantly constrains generative models’ de novo design capabilities. While incorporating 1D sequence information improved D-Flow’s performance by 0.58%, this impact was less pronounced than in previous protein-related tasks [77] like ligand docking [14], ligand efficacy prediction [66], or model quality assessment [13]. We attribute this to PLMs’ inherent limitations with short sequences like peptides, compared to their effectiveness with longer proteins. This gap generally arises because PLMs learn contextual relationships by observing long-range dependencies and conserved motifs – key elements that characterize protein families and evolutionary relationships [56, 21]. In short sequences, like peptides, these relationships are limited due to a lack of length and structure, making it harder for the model to capture evolutionary patterns effectively. Peptides often lack secondary and tertiary structure information present in longer sequences, which is critical for understanding function and evolutionary context [4]. Those attention-based models, such as transformers [68], distribute attention weights over the input length. Shorter sequences provide fewer tokens to which attention can be allocated, which can limit the richness of the learned representation. Regarding other components, DFM demonstrated superior performance over the probability simplex mechanism [41] for amino acid type representation, improving AAR by 1.43%. The control technique provided a modest 0.32% enhancement in AAR.
[…]
1 - - - - 50.43
2 - - - 51.86 +1.43
3 - - 52.44 +0.58
4 - 58.37 +5.93
5 58.69 +0.32
Table 2: Ablation studies on the effects of each module.
Figure 2: Comparison of L- and D-peptide examples generated by D-Flow with the same target receptor, where the red one is the ground truth L-peptide.
5.2 D-peptide Design
Generation of Pure D-peptides.
@@ 5.2 D-peptide Design
Our analysis confirms that D-Flow successfully produces pure D-peptides in all cases (Fig. 3). Remarkably, this achievement occurs without any D-protein training corpus, requiring only our specialized post-processing technique $\Psi(\cdot)$. To understand it, as D-Flow gradually moves the peptide’s sequence and structure $(a,\boldsymbol{x},O,\boldsymbol{\chi})$ from a prior distribution $p_{0}$ to the target data distribution $p_{1}$ at a speed of $v_{t}$ during $t\in[0,1]$, its moving trajectory is heavily dependent on the anchor fixed receptor $\mathcal{C}_{\text{rec}}$. The velocity vectors $v^{\text{pos}}(\cdot),v^{\text{ori}}(\cdot)$, and $v^{\text{ang}}(\cdot)$ for $\mathrm{C}\alpha$ locations, frame orientations, and angles are computed based on the interactions between the intermediate state of peptide $\mathcal{C}_{\text{pep}}^{t}$ and the target protein $\mathcal{C}_{\text{rec}}$. Once a D-chirality receptor $\Psi(\mathcal{C}_{\text{rec}})$ is provided as the condition, our D-Flow $f_{\theta}(\cdot)$ captures the chirality nuance of this given input. As a response, it adjusts $v^{\text{pos}}(\cdot),v^{\text{ori}}(\cdot)$, and $v^{\text{ang}}(\cdot)$ to accord with this D-target protein, which leads to a product with considerable D-residues.
Model Robustness Assessment.
@@
We examined the robustness of D-Flow by testing multiple mirror-image approaches for the input receptor. Despite there being infinite possible mirror transformations, we focused on the three fundamental axes (x, y, and z). Our results demonstrate that the chirality of generated peptides remains consistently D-configured regardless of the mirroring approach used.
Conformational Distribution Analysis
@@
Fig. 2 visualizes a few examples of designed L- and D-peptides. While D-Flow successfully generates both configurations, our analysis reveals a striking contrast: L-peptides show remarkable consistency in their structural and sequence distributions, whereas D-peptides exhibit significantly higher variability, particularly in longer sequences. This phenomenon can be attributed to several fundamental factors.
@@
Training Distribution Mismatch. Our model, trained exclusively on L-peptides, has learned to navigate L-peptide conformations’ natural distribution and energy landscape. When presented with a D-receptor through the mirror-image algorithm, the model must operate in a conformational space absent from its training distribution. This distributional shift forces the model to explore the D-space with less confidence, increasing structural diversity [65]. Loss of Evolutionary Constraints. L-peptides in our training data reflect millions of years of evolutionary optimization, containing implicit biases about energetically favorable conformations [23, 40]. While the mirror-image transformation preserves theoretical symmetry, these evolutionary constraints don’t translate perfectly to the D-space, potentially contributing to the observed variability. Asymmetric Energy Landscape Exploration. The FM process relies on smooth transitions across the conformational space [44]. However, the mirror-image transformation introduces subtle numerical asymmetries in the computational representation of D-peptides. These asymmetries can affect how the model explores the energy landscape, leading to more diverse intermediate states and final structures [74]. Stereochemical Environment Adaptation. The significant change in the stereochemical environment during mirroring poses a unique challenge. The flow model must adapt its learned representations of molecular interactions to account for the inverted chirality [5]. This adaptation process isn’t perfect, resulting in broader conformational sampling and increased structural diversity in the generated D-peptides.
@@
These challenges align with recent findings in protein design that highlight the challenges of transferring learned molecular representations across different stereochemical spaces [62]. The energy landscape learned for L-peptides doesn’t directly translate to the D-space due to fundamental differences in molecular interactions and packing arrangements. This leads to broader structural exploration and increased diversity in the generated structures [52]. This analysis suggests potential avenues for improvement, such as incorporating symmetry-aware constraints or developing specialized adaptation mechanisms for D-peptide generation, which could help stabilize the output distribution while maintaining the desired chirality.
Figure 3: Perspective view of D-residues within the D-peptides, where stereogenic carbon alpha to the amino group has the D-configuration.

### v3

@@ 2.2 Parameterization with Adapter-guided Protein Language Models
To efficiently learn $(a^{i},\boldsymbol{x}^{i},O^{i},\boldsymbol{\chi}^{i})$ for every residue, we adopt and improve the FramePred architecture [67, 66]. It incorporates Invariant Point Attention (IPA) [18] to encode spatial features and ensure equivariance. IPA consists of multiple Transformer layers [50] to capture sequence-level attributes. In addition, considering the periodicity, torsion angles $\boldsymbol{\chi}\in[0,2\pi)^{7}$ are flexibly encoded by applying multi-frequency sine and cosine transformations [26], which are also fused with the timestep embedding and residue sequence embedding into IPAs.
PLMs capture the evolutionary patterns from large-scale sequence data, and this knowledge is proven to be supportive of protein folding [28] and inverse design [71], which is evidently beneficial for our co-design task as well. Some prior studies [62] integrate this knowledge by immediately appending the geometric networks to PLMs. Drawing inspiration from LM-Design [71], we employ a lightweight structural adapter to endow PLMs with structural awareness. In our approach, if the IPA output is denoted as $\boldsymbol{h}_{\text{IPA}}\in\mathbb{R}^{\psi_{\text{IPA}}}$, the $l$-th layer’s attention is computed as: $\boldsymbol{o}=\mathrm{softmax}\left(\frac{\boldsymbol{h}_{\text{seq}}\boldsymbol{W}_{Q}\cdot\boldsymbol{h}_{\text{IPA}}^{\top}\boldsymbol{W}_{K}^{\top}}{\sqrt{\psi_{\text{seq}}}}\right)\boldsymbol{h}_{\text{IPA}}\boldsymbol{W}_{V}$, where $\boldsymbol{h}_{\text{seq}}\in\mathbb{R}^{\psi_{\text{seq}}}$ is the sequential embedding of the last (l1)(l-1)-th layer. $\boldsymbol{W}_{K}$, $\boldsymbol{W}_{Q}$, and $\boldsymbol{W}_{V}$ are trainable weights for key, query, and values, respectively. This module allows the usage of sequence- and structure-based information for protein understanding.
@@ 2.3 Controlling Flow Matching with Target Proteins
Existing de novo design algorithms [66, 67] are mainly targeting monomers or biomolecules without receptors. Some recent studies [31, 26, 27] have considered binder design but their performance is strictly restricted by the number of available complex structures (e.g., SabDAB [13] and PepMerge [26]). To bridge the gap and exploit all crystal structures, we pretrain our flow model on the vast amount of general proteins (i.e., Protein Data Bank) and then fine-tune it on the target-specific design challenges.
To begin with, $\mathcal{F}_{\Theta}:\mathcal{P}_{0}\rightarrow\mathcal{C}_{\text{pro}}$ is a co-design flow model without target awareness, which can de novo design a protein ${C}_{\text{pro}}$ from any initial protein ${C}_{0}$ that is drawn from a prior distribution $p_{0}$. It is trained on general proteins to approximate $p(a,\boldsymbol{x},O,\boldsymbol{\chi})$. To enable awareness of $\mathcal{F}_{\Theta}$ to the receptor $C_{\text{rec}}$, we borrow ideology from ControlNet [69], locking (freeze) the parameters $\Theta$ of the original block and simultaneously cloning the block to a trainable copy with parameters $\Theta_{c}$. The trainable copy takes a receptor protein $C_{\text{rec}}$ as input and is connected to the locked model with zero IPAs denoted as $\mathcal{Z}(\cdot)$. In particular, both the weight and the bias in $\mathcal{Z}(\cdot)$ are initialized to zeros. In practice, two instances of zero IPAs are used with parameters $\Theta_{\mathrm{z}1}$ and $\Theta_{\mathrm{z}2}$, respectively:
@@
2.4 D-peptide Design
In this work, we generate mini D-protein binders, specifically D-peptides, using the mirror-image algorithm [45]. The mirror-image operation, denoted as $\Psi(\cdot)$, transforms any given protein structure into its corresponding mirror image. $\Psi(\cdot)$ inverts the spatial configuration of the protein, effectively reflecting it across an imaginary plane while preserving the relative distances and angles between atoms. In our approach, we first apply $\Psi(\cdot)$ to the receptor protein $C_{\text{rec}}$, converting it into its mirror-image counterpart. Then, we pass this mirrored receptor through a flow model $f_{\theta}(\cdot):\mathcal{C}_{\text{rec}}\rightarrow\mathcal{C}_{\text{pep}}$, which is trained on natural-chirality (L-) proteins, to yield the corresponding peptide. Finally, we apply the inverse operation $\Psi^{-1}(\cdot)$ to convert the resulting peptide back to the D-configuration. Formally, the generation process can be expressed as:
$C^{\text{D-pep}}=\Psi^{-1}(f_{\theta}(\Psi(C_{\text{rec}}))).$ (16)
Notably, this reflection $\Psi(\cdot)$ can be represented by a parity transformation, written as $\Psi(C)=MC$. where $M$ is a reflection matrix and satisfies $M^{T}M=I$ and $\operatorname{det}(M)=-1$. Unlike rotation matrices (which have determinant +1 and preserve orientation), a reflection matrix $M$ flips orientation and reverses chirality (e.g., converting an L-protein to a D-protein). Since mirroring preserves pairwise distances and angles but reverses chirality, it commutes with translations but interacts with rotations depending on their handedness. Therefore, the following conclusion holds:
Theorem 1
If $f_{\theta}$ is roto-translational equivariant, then the mirror-image algorithm in Equ. 16 maintains the equivariance property. That is $\Psi^{-1}\left(f_{\theta}\left(\Psi\left(RC_{\mathrm{rec}}+\text{t}\right)\right)\right)=R\Psi^{-1}\left(f_{\theta}\left(\Psi\left(C_{\mathrm{rec}}\right)\right)\right)+\text{t}.$, where $(R,\text{t})$ is the rigid transformation.
It indicates that applying any rigid transformation (rotation $R$ and translation $t$) to the input receptor structure $C_{\text{rec}}$ before running it through the mirror-image algorithm results in the same transformation applied to the output $C^{\text{D-pep}}$. The proof of the theorem is in C.
3 Results
We execute two types of experiments to validate the effectiveness of our D-Flow approach. The first, detailed in Sec. 3.1, involves the conventional co-design challenge for L-proteins, where models generate peptides’ both sequences and structures based on a given receptor binding site. The second experiment, described in Sec. 3.2, focuses on generating D-proteins. For benchmarking, we use the PepMerge dataset [26], derived from PepBDB [54] and Q-BioLip [53]. To ensure a fair comparison with prior work [26], we cluster peptide-protein complexes based on 40% sequence identity using MMseqs2 [43], after removing duplicates and applying empirical filters (e.g., resolution << 4Å, peptide length between 3 and 25). This process resulted in 8,365 non-redundant complexes across 292 clusters. For consistent comparison, we use the same test set as Li and others [26], containing 10 clusters and 158 complexes. Pretraining data contain monomers between lengths 60 and 384 with resolution $5\textup{\AA }$ from PDB [3] on August 8, 2021, ensuring no data leakage for peptides. The data is then filtered by including proteins with high secondary structure compositions only. Monomers with more than 50% loops are also removed using DSSP [19], resulting in 18,684 proteins.
3.1 Unconditioned Sequence-structure Co-design
Baselines. We select two categories of state-of-the-art protein design methods as baselines. The first category disregards side-chain conformations and includes approaches such as RFDiffusion [52] and ProteinGen [30]. RFDiffusion generates protein backbones, with sequences predicted afterward using ProteinMPNN [8], while ProteinGen improves on RFDiffusion by jointly sampling both backbones and sequences. The second category considers full-atom protein generation, containing Diffusion [31], PPIFlow [27], and PepFlow [26]. PepFlow has three variants based on whether backbones, sequences, and side-chain angles are sampled. For each test case, we sample 64 peptides simultaneously.
Table 1: Evaluation of methods in the traditional L-peptide sequence-structure co-design task, where metrics are divided into three main categories. The best and suboptimal results are labeled boldly and underlined, respectively.
@@ 3.1 Unconditioned Sequence-structure Co-design
AAR $\%\uparrow$ RMSD $\text{\AA }\downarrow$ $\operatorname{SSR}\%\uparrow$ BSR $\%\uparrow$ Stb. $\%\uparrow$ Aff. $\%\uparrow$ Des. $\%\uparrow$ Div. $\uparrow$
@@
Table 2: Ablation studies on the effects of each module. To be specific, DFM is the abbreviation of discrete flow matching, PLM is the usage of protein language models, Pretrain is the transfer learning from PDB to PepMerge, and CN is the employment of the ControlNet technique.
[…]
1 - - - - 50.43 2.34
2 - - - 51.86 +1.43 2.09 - 0.25
3 - - 52.44 +0.58 2.01 -0.08
4 - 58.37 +5.93 1.78 -0.23
5 58.69 +0.32 1.63 -0.15
Main Results. Generated peptides are evaluated across three key aspects:Geometry, Energy, and Design (Detailed see Appendix D).As documented in Tab. 1, D-Flow generates peptide sequences with the closest resemblance to native ones, achieving a high amino acid identity of 58.69%, a 14.51% improvement over PepFlow. It excels in all geometry-related metrics, highlighting its close alignment with the binding sites of native peptides. Besides, D-Flow scores 75.14% in designability and 0.60 in diversity, maintaining a good balance between structural fidelity and variety. Moreover, D-Flow also demonstrates superior energy-based properties, achieving the best stability score at 26.85% and the best affinity score at 24.31%, which are critical for the formation of strong and stable complexes. In a nutshell, D-Flow outperforms all baselines across most key metrics, indicating its strength in consistently producing the most accurate peptide sequences and structures, along with the optimal stability and affinity for their targets.
Figure 2: Perspective view of D-residues within two randomly selected D-peptides generated by our D-flow, where all stereogenic alpha carbons to the amino group in D-peptides have the D-configuration.
We also navigate the contribution of D-Flow components, containing the discrete flow matching (DFM), structural-adapted PLM, the pretraining on PDB, and the ControlNet-style transfer learning technique. Tab. 2 reports the ablation study results. It can be observed that the integration of additional unlabeled structural data yielded the most substantial improvement, increasing AAR by 5.93%. This supports our hypothesis that limited binding complex data significantly constrains generative models’ de novo design capabilities. While incorporating 1D sequence information improved D-Flow’s performance by 0.58%, this impact was less pronounced than in previous protein-related tasks [62, 61] like ligand docking [17, 23], ligand efficacy prediction [10], or model quality assessment. We attribute this to PLMs’ inherent limitations with short sequences like peptides, compared to their effectiveness with longer proteins. This gap generally arises because PLMs learn contextual relationships by observing long-range dependencies and conserved motifs – key elements that characterize protein families and evolutionary relationships [39]. In short sequences, such as peptides, these relationships are limited due to their short length and lack of structure, making it harder for PLMs to capture evolutionary patterns effectively. Peptides often lack secondary and tertiary structure information present in longer sequences, which is critical for understanding function and evolutionary context. Attention-based models like transformers [50] distribute attention weights over the input length. Shorter sequences provide fewer tokens to which attention can be allocated, limiting the richness of learned representations. Regarding other components, DFM demonstrated superior performance over the probability simplex mechanism [26] for amino acid type representation, improving AAR by 1.43%. The control technique provided a modest 0.32% enhancement in AAR.
Figure 3: Visualization of L- and D-peptide examples generated by D-Flow with the same target receptor, where the red one is the ground truth L-peptide. Despite that D-flow achieves particularly high recovery of both amino acid sequences and structures for L-peptides, its sampling space of D-peptides is relatively diverse and unconstrained. This phenomenon of diversity can be explained by the distribution mismatch between the training and inference data, loss of evolutionary constraints, asymmetric energy landscape exploration, and stereochemical environment adaptation.
@@ 3.2 D-peptide Design
Generation of Pure D-peptides. Our analysis confirms that D-Flow successfully produces pure D-peptides in all cases (Fig. 2). Remarkably, this achievement occurs without any D-protein training corpus, requiring only our specialized post-processing technique $\Psi(\cdot)$. To understand it, as D-Flow gradually moves the peptide’s sequence and structure $(a,\boldsymbol{x},O,\boldsymbol{\chi})$ from a prior distribution $p_{0}$ to the target data distribution $p_{1}$ at a speed of $v_{t}$ during $t\in[0,1]$, its moving trajectory is heavily dependent on the anchor fixed receptor $\mathcal{C}_{\text{rec}}$. The velocity vectors $v^{\text{pos}}(\cdot),v^{\text{ori}}(\cdot)$, and $v^{\text{ang}}(\cdot)$ for $\mathrm{C}\alpha$ locations, frame orientations, and angles are computed based on the interactions between the intermediate state of peptide $\mathcal{C}_{\text{pep}}^{t}$ and the target protein $\mathcal{C}_{\text{rec}}$. Once a D-chirality receptor $\Psi(\mathcal{C}_{\text{rec}})$ is provided as the condition, our D-Flow $f_{\theta}(\cdot)$ captures the chirality nuance of this given input. As a response, it adjusts $v^{\text{pos}}(\cdot),v^{\text{ori}}(\cdot)$, and $v^{\text{ang}}(\cdot)$ to accord with this D-target protein, which leads to a product with considerable D-residues.
@@
Model Robustness Assessment. We examined the robustness of D-Flow by testing multiple mirror-image approaches for the input receptor. Despite there being infinite possible mirror transformations, we focused on the three fundamental axes (x, y, and z). Our results demonstrate that the chirality of generated peptides remains consistently D-configured regardless of the mirroring approach used.
@@
Conformational Distribution Analysis. Fig. 3 visualizes a few examples of designed L- and D-peptides. While D-Flow successfully generates both configurations, our analysis reveals a striking contrast: L-peptides show remarkable consistency in their structural and sequence distributions, whereas D-peptides exhibit significantly higher variability, particularly in longer sequences. This phenomenon can be attributed to several fundamental factors.
@@
Training Distribution Mismatch. Our model, trained exclusively on L-peptides, has learned to navigate L-peptide conformations’ natural distribution and energy landscape. When presented with a D-receptor through the mirror-image algorithm, the model must operate in a conformational space absent from its training distribution. This distributional shift forces the model to explore the D-space with less confidence, increasing structural diversity. Loss of Evolutionary Constraints. L-peptides in our training data reflect millions of years of evolutionary optimization, containing implicit biases about energetically favorable conformations [25]. While the mirror-image transformation preserves theoretical symmetry, these evolutionary constraints don’t translate perfectly to the D-space, potentially contributing to the observed variability. Asymmetric Energy Landscape Exploration. The FM process relies on smooth transitions across the conformational space [29]. However, the mirror-image transformation introduces subtle numerical asymmetries in the computational representation of D-peptides. These asymmetries can affect how the model explores the energy landscape, leading to more diverse intermediate states and final structures [55]. Stereochemical Environment Adaptation. The significant change in the stereochemical environment during mirroring poses a unique challenge. The flow model must adapt its learned representations of molecular interactions to account for the inverted chirality [2]. This adaptation process isn’t perfect, resulting in broader conformational sampling and increased structural diversity in the generated D-peptides.
@@
These challenges align with recent findings in protein design that highlight the challenges of transferring learned molecular representations across different stereochemical spaces [44]. The energy landscape learned for L-peptides doesn’t directly translate to the D-space due to fundamental differences in molecular interactions and packing arrangements. This leads to broader structural exploration and increased diversity in the generated structures. This analysis suggests potential avenues for improvement, such as incorporating symmetry-aware constraints or developing specialized adaptation mechanisms for D-peptide generation, which could help stabilize the output distribution while maintaining the desired chirality.